OpenStudy (anonymous):

So i have a math page that i need some answers for. 1/2 * 3/5 3/8* 16 7/15*10/21 1 1/3* 1 1/5 25* 2 1/5 They need to be in simplest form

OpenStudy (amistre64):

multiply straight across; then reduce 1 3 --- = 3/10 ; cant reduce so its in simplest form 2 5

OpenStudy (amistre64):

3.16 ---- can be reduced now or later, but is usually easier to reduce now 8 3.2 --- = 6 1

OpenStudy (amistre64):

7.10 ----- can reduce now 15.21 1.2 --- = 2/9 ; simplest form 3.3

OpenStudy (amistre64):

1' 1/3 . 1' 1/5 should change to improper fractions and do the same

OpenStudy (amistre64):

4.6 --- reduce now 3.5 4.2 --- = 8/5 keep as is or revert to mixed numerals form of: 1' 3/5 5